# Pilot-waves and copilot-particles: A novel approach to objective collapse

### v2

@@ Abstract
We propose an extension of Schrödinger’s equation that incorporates the macroscopic measurement-induced wavefunction collapse phenomenon. Our approach relies on a hybrid between Bohm-de Broglie pilot-wave and objective collapse theories. The Bohmian particle is guided by the wavefunction and, conversely, the wavefunction gradually localizes towards the particle’s position. As long as the particle can visit any state, as in a typical microscopic system, the localization effect does not favor any particular quantum state and, on average, the usual Schrödinger-like time evolution results. However, when the wavefunction develops spatially well-separated lobes, as would happen during a macroscopic measurement, the Bohmian particle can remain trapped in one lobe, and the wavefunction eventually localizes there. The end result, in macroscopic systems, is a wavefunction collapse that is consistent with Born’s rule. We illustrate the theory with a simple double-slit experiment simulation.
@@
The theory is more naturally presented in the context of the Schrödinger equation for continuous degrees of freedom (see End Matter for a treatment of discrete states). Below, $\psi\left(x,t\right)$ is a properly (anti-)symmetrized wavefunction in position representation, with $t$ being time and $x$ being a $3N$-dimensional vector for a system of $N$ particles. (We consider the familiar context of “first quantization”, but there is no fundamental problem associated with applying it to boson fields within quantum field theory [20], in which case $x$ would represent a specific value of a classical field.) Like most theories with either a pilot-wave or an objective collapse flavor, our theory is non-relativistic. Our proposed modified Schrödinger equation takes the form:
@@
There remains to specify to equation of motion for $r$. We depart from the standard Bohm-de Broglie prescription and instead assume that the particle experiences a fictitious potential $-\tau\ln\left|\psi\left(r,t\right)\right|^{2}$, where $\tau$ is a constant, and otherwise obeys a classical equation of motion:
@@
where $\mu$ is a $3N\times 3N$ diagonal matrix of fictitious masses. We allow for a time delay $\Delta t$ in the effect of the wavefunction (having $\left|\psi\left(r,t-\Delta t\right)\right|$ instead of the instantaneous effect $\left|\psi\left(r,t\right)\right|$) to avoid a self-interaction artifact (to be discussed in more detail below). Standard statistical mechanics arguments [21] imply that, in equilibrium at a fictitious temperature $\tau$ (expressed in energy units), the position $r$ will be distributed as $\exp\left(-\left(-\tau\ln\left|\psi\left(r,t\right)\right|^{2}\right)/\tau\right)=\left|\psi\left(r,t\right)\right|^{2}$, as desired. This presupposes that the velocities are initialized (in the distant past) consistently with the fictitious temperature $\tau$, e.g. $\frac{dr_{k}}{dt}\sim\operatorname{Normal}\left(0,\mu_{k}^{-1}\tau\right)$ for $k=1,\ldots,3N$. This argument also assumes that the particle equilibrates rapidly relative to the time scale of changes in the wave function magnitude $\left|\psi\left(r,t\right)\right|$. Note that all “fictitious” quantities are not tied to any real observable properties of the system: $\mu$ is not necessarily related to the actual particles’ masses and $\tau$ may not be the real temperature. In fact, these quantities enter all equations as the ratio $\mu^{-1}\tau$, which we leave as an adjustable parameter in the theory. It is interesting that the logarithm in Equation (2) coincidentally accomplishes two tasks: separating out the magnitude from the phase (since $\ln\left|\psi\left(r,t\right)\right|^{2}=2\operatorname{Re}\ln\psi\left(r,t\right))$ and placing the magnitude on a logarithmic scale, as needed to cancel out the Boltzmann factor.
@@
Figure 1: (Color online) Microscopic vs. macroscopic behaviors. Upper left: For a spatially localized wavefunction, the Bohmian particle (represented by a grey spheres trail) can access all regions where the wavefunction is nonnegligible (here represented by fictitious potential surface proportional to $-\ln|\psi(r,t)|^{2}$). Middle: As a superposition of states starts to involve spatially well-separated states, the particle remains trapped in one of the “lobes” of the wavefunction. Lower right: Loss of ergodicity causes wavefunction localization around one of the macroscopically distinct outcomes
@@
The reason we deviate from standard Bohm-de Broglie dynamics are twofold. First, the particles tend to not get trapped very well under the usual Bohm-de Broglie dynamics: As the particle approaches a region of low probability density, it typically speeds up (which specifically ensures that the particle has low probability of being found there). As a result, the particle can easily escape any isolated lobe in the wavefunction. Second, the Bohmian particle does not thermalize very quickly. For instance, in a wave packet, the particle typically follows along smoothly while remaining at the same location relative to the center of the wave packet, instead of visiting all positions of high probability. Our alternative dynamics avoids both of these problems.
@@
We are now ready to show that, under conditions ensuring thermalization, our proposal essentially reduces to the conventional Schrödinger equation. Indeed, computing the expected value of the last term of (1) with respect to the random position $r$ with instantaneous probability density $\left|\psi\left(r,t\right)\right|^{2}$, we have:
@@
[…]
In contrast, as illustrated in Figure 1, if the wavefunction develops spatially well-separated lobes and the particle is trapped in only one of the them, then the last term remains essentially zero for the lobes that are no longer visited and, there, only the decaying term $-\kappa\psi\left(x,t\right)$ is active, resulting in an exponential decay of the wavefunction. Within the lobe trapping the particle, the localization term of (1) is larger than average, since the particle visit points there in more often than would be indicated by the wavefunction magnitude. The end result is that the localization terms there exceeds the decaying term on average and the wavefunction gradually concentrates onto that lobe at the expense of the others. The loss of ergodicity of the Bohmian particle is the source of the wavefunction collapse within this model. This helps explain single outcomes in macroscopic measurements, since they arguably lead to wavefunctions with macroscopically disconnected lobes, such as a dial (or “pointer”) being in one place or another or a photon being emitted in different directions, etc. The End Matter extends this reasoning to the measurement of a subsystem.
Similarly, another mechanism that could cause a loss of ergodicity is the process of copying a microscopic state onto a large number of particles (within the limits of the no-cloning theorem). This operation typically results in increasingly deeper basins in the fictitious potential $-\tau\ln\left|\psi\left(r,t\right)\right|^{2}$ that increase the likelihood of the particle being trapped. To see this, consider a particle in a superposition of two states $\alpha\left|1\right\rangle+\beta\left|2\right\rangle$ (with $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$) that is starting to interact with a measurement apparatus made of $N-1$ distinguishable particles, initially in a “no measurement” state $\left|0\right\rangle\otimes\cdots\otimes\left|0\right\rangle$. After a suitable unitary evolution, the joint system transitions to the state $\alpha\left|1\right\rangle\otimes\left|1\right\rangle\otimes\cdots\otimes\left|1\right\rangle+\beta\left|2\right\rangle\otimes\left|2\right\rangle\otimes\cdots\otimes\left|2\right\rangle$. Now assume that, for particle $i$, the states $\left|1\right\rangle$ and $\left|2\right\rangle$ are, respectively, associated with single-particle wavefunctions in position representation $\psi_{1}\left(x_{i}\right)=\varphi\left(x_{i}-a\right)$ and $\psi_{2}\left(x_{i}\right)=\varphi\left(x_{i}+a\right)$ for some macroscopic value $a\in\mathbb{R}^{3}$ and centrosymmetric unimodal function $\varphi\left(x_{i}\right)$ (with $x_{i}\in\mathbb{R}^{3}$). The overall wavefunction is then
Figure 2 illustrates our proposed time evolution of the wavefunction in a classic setting: the double-slit experiment. The simulation shows both the initial wave-like interference behavior and the eventual collapse to a more specific outcome as the wave packet heads towards a detector. The Supplementary materials include an animation of the same numerical experiment, as well as the simulation code itself. The End Matter also includes a treatment of a standard EPR-type experiment [22].
[…]
@@
The idea that the wavefunction is attracted to particles can be found in earlier work [18], but our approach differs in substantial ways. We consider collapse towards a point in $\mathbb{R}^{3N}$ instead of a towards a single-particle density. Keeping $3N$-dimensional information helps maintain EPR-type correlations between distant measurements. Also, unlike [18], we deviate from standard Bohmian dynamics to ensure particle trapping.
@@
The proposed theory is compatible with the standard decoherence picture (see End Matter for a density matrix formulation): it does not affect the basic idea that interactions with the environment causes a decay of the off-diagonal terms of a subsystem’s density matrix. But once the resulting mixture of outcomes involves macroscopically distinct states, our equation of motion predicts localization towards one of these macroscopic states. Hence, it provides an objective “selection rule” that is absent under decoherence alone and avoids the need for a many-world interpretation. The decoherence concept makes it unnecessary for the wave function to collapse rapidly in our framework. The conversion from a superposition of states to a statistical mixture takes place over the same time span under our theory as under decoherence alone. The exponential decay of all but one branch of the statistical mixture, however, can take much longer: any decay rate faster than human perception (i.e., milliseconds) would be sufficient to give human observers the experience of a single measurement outcome.
The proposed theory does exhibit some undesirable features (see End Matter). Most obviously, the proposed modified Schrödinger equation is slightly nonlinear, whereas linearity of quantum mechanics has been verified experimentally to a great precision [23]. This places a tight upper bound on the parameter $\kappa$, thus also placing a bound on the speed at which wavefunction collapse can occur. The fact that the nonlinearity must be small mitigates other known issues associated with nonlinearity [24].
@@
While the current treatment is non-relativistic, nothing seems to forbid a relativistic extension. The collapse phenomenon involves local operations: the decay term $-\kappa\psi\left(x,t\right)$ is purely local while the localization terms only involves instantaneous effects over short distances (as governed by the function $\phi_{r}\left(x\right)$) and ensuring that this effect propagates at a finite speed entails no significant changes in the theory. Since relativistic versions of pilot-wave [25] and objective collapse theories [26] have been developed with some success, similar ideas could carry over in the present setup.

### v3

# Pilot-waves and copilot-particles: A nonstochastic approach to objective wavefunction collapse

@@ Abstract
We seek an extension to Schrödinger’s equation that incorporates the macroscopic measurement-induced wavefunction collapse phenomenon. We find that a suitable hybrid between two leading approaches, the Bohm-de Broglie pilot-wave and objective collapse theories, accomplishes this goal in a way that is consistent with Born’s rule. Our theory posits that the Bohmian particle is guided by the wavefunction and, conversely, the wavefunction gradually localizes towards the particle’s position. As long as the particle can visit any state, as in a typical microscopic system, the localization effect does not favor any particular quantum state and, on average, the usual Schrödinger-like time evolution results. However, when the wavefunction develops spatially well-separated lobes, as would happen during a macroscopic measurement, the Bohmian particle can remain trapped in one lobe, which causes the wavefunction to eventually localizes. This proposed loss of ergodicity mechanism recasts one of the foundational postulate of quantum mechanics as a emergent feature and has important implications regarding the feasibility of large-scale quantum computing.
[…]
Measurement postulate, Bohmian mechanics, Nonlinear wave equation
1 Introduction
@@
2 Theory
The theory is more naturally presented in the context of the Schrödinger equation for continuous degrees of freedom (see Appendix E for a treatment of discrete states). Below, $\psi\left(x,t\right)$ is a properly (anti-)symmetrized wavefunction in position representation, with $t$ being time and $x$ being a $3N$-dimensional vector for a system of $N$ particles. (We consider the familiar context of “first quantization”, but there is no fundamental problem associated with applying it to boson fields within quantum field theory [20], in which case $x$ would represent a specific value of a classical field.) Like most theories with either a pilot-wave or an objective collapse flavor, our theory is non-relativistic. Our proposed modified Schrödinger equation takes the form:
@@ 2 Theory
There remains to specify to equation of motion for $r$. We depart from the standard Bohm-de Broglie prescription and instead assume that the particle experiences a fictitious potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$, where $\tau$ is a constant, and otherwise obeys a classical equation of motion:
@@
where $\mu$ is a $3N\times 3N$ diagonal matrix of fictitious masses. We allow for a time delay $\Delta t$ in the effect of the wavefunction (having $\left|\psi\left(r,t-\Delta t\right)\right|$ instead of the instantaneous effect $\left|\psi\left(r,t\right)\right|$) to avoid a self-interaction artifact (to be discussed in more detail below). Standard statistical mechanics arguments [21] imply that, in equilibrium at a fictitious temperature $\tau$ (expressed in energy units), the position $r$ will be distributed as $\exp\left(-\left(-\tau\ln\left|\psi\left(r,t\right)\right|^{2}\right)/\tau\right)=\left|\psi\left(r,t\right)\right|^{2}$, as desired. This presupposes that the velocities are initialized (in the distant past) consistently with the fictitious temperature $\tau$, e.g. $\frac{dr_{k}}{dt}\sim\operatorname{Normal}\left(0,\mu_{k}^{-1}\tau\right)$ for $k=1,\ldots,3N$. The velocity degrees of freedom can play the role of thermal bath, so no additional thermal bath is needed. This argument also assumes that the particle equilibrates rapidly relative to the time scale of changes in the wave function magnitude $\left|\psi\left(r,t\right)\right|$. This equilibration behavior is implied by Boltzmann’s classic “H-theorem”, although it is difficult to formulate simple general conditions guaranteeing a given rate of convergence (for a recent account of these types of results, see [22]). In fact, in the sequel we assume that such fast equilibration behavior is guaranteed only for microscopic systems. A simple dimensional argument indicates that the equilibration time should scale with $\ell/\sqrt{\mu^{-1}\tau}$, where $\ell$ is of the order of magnitude of the linear spatial extent probability distribution of a particle’s position.
Note that all quantities labelled here as “fictitious” are not tied to any real observable properties of the system: $\mu$ is not necessarily related to the actual particles’ masses and $\tau$ may not be the real temperature. In fact, these quantities enter all equations as the ratio $\mu^{-1}\tau$, which we leave as an adjustable parameter in the theory that controls the equilibration time. It is interesting that the logarithm in Equation (2) coincidentally accomplishes two tasks: separating out the magnitude from the phase (since $\ln\left|\psi\left(r,t\right)\right|^{2}=2\operatorname{Re}\ln\psi\left(r,t\right))$ and placing the magnitude on a logarithmic scale, as needed to cancel out the Boltzmann factor.
[…]
@@
Figure 1: (Color online) Microscopic vs. macroscopic behaviors. Upper left: For a spatially localized wavefunction, the Bohmian particle (represented by a grey spheres trail) can access all regions where the wavefunction is nonnegligible (here represented by fictitious potential surface proportional to $\ln|\psi(r,t)|^{2}$). Middle: As a superposition of states starts to involve spatially well-separated states, the particle remains trapped in one of the “lobes” of the wavefunction. Lower right: Loss of ergodicity causes wavefunction localization around one of the macroscopically distinct outcomes
@@
Traditional Bohm-de Broglie dynamics guarantees a strong equivariance property: If an ensemble of Bohmian particles is distributed according the $\left|\psi\left(r,t_{0}\right)\right|^{2}$ at some time $t_{0}$, it will remain, by construction, distributed according to $\left|\psi\left(r,t\right)\right|^{2}$ for all $t>t_{0}$. However, we deviate from standard Bohm-de Broglie dynamics for two reasons. First, the particles tend to not get trapped very well under the usual Bohm-de Broglie dynamics: As the particle approaches a region of low probability density, it typically speeds up (which specifically ensures that the particle has low probability of being found there). As a result, the particle can easily escape any isolated lobe in the wavefunction. Second, the Bohmian particle does not thermalize very quickly. For instance, in a wave packet, the particle typically follows along smoothly while remaining at the same location relative to the center of the wave packet, instead of visiting all positions of high probability. Our alternative dynamics avoids both of these problems. The equivariance property under Bohm-de Broglie dynamics is here replaced by a convergence towards $\left|\psi\left(r,t\right)\right|^{2}$ implied by Boltzmann’s “H-theorem” [22]. When the relaxation time associated with this process become very long, the particle’s distribution could differ from $\left|\psi\left(r,t\right)\right|^{2}$ and this possibility will in fact turn out to be essential to enable wavefunction collapse in macroscopic systems in our formalism.
@@
We are now ready to show that, under conditions ensuring rapid thermalization (i.e. when the particle’s relaxation time is sufficiently short), our proposal essentially reduces to the conventional Schrödinger equation. Indeed, computing the expected value of the last term of (1) with respect to the random position $r$ with instantaneous probability density $\left|\psi\left(r,t\right)\right|^{2}$, we have:
$\int_{r\in\mathbb{R}^{3N}}\kappa\bar{\phi}_{r}\left(x\right)\frac{1}{\left|\psi\left(r,t\right)\right|^{2}}\left|\psi\left(r,t\right)\right|^{2}\psi\left(x,t\right)\penalty 10000\ dr$
$=\kappa\frac{1}{\#\Pi}\sum_{\pi\in\Pi}\int_{r\in\mathbb{R}^{3N}}\phi\left(x-\pi r\right)dr\penalty 10000\ \psi\left(x,t\right)=\kappa\psi\left(x,t\right)$ (3)
since $\int\phi\left(x-\pi r\right)dr=1$ by construction. As a result, on average, the last term cancels out the second term of (1) and we recover Schrödinger’s equation.
In contrast, as illustrated in Figure 1, if the wavefunction develops spatially well-separated lobes and the particle is trapped in only one of the them (i.e. when the particle’s global relaxation time becomes very long), then the last term remains essentially zero for the lobes that are no longer visited and, there, only the decaying term $\kappa\psi\left(x,t\right)$ is active, resulting in an exponential decay of the wavefunction. Within the lobe trapping the particle, the localization term of (1) is larger than average, since the particle visit points there in more often than would be indicated by the wavefunction magnitude. The end result is that the localization terms there exceeds the decaying term on average and the wavefunction gradually concentrates onto that lobe at the expense of the others. The loss of ergodicity of the Bohmian particle is the source of the wavefunction collapse within this model. This helps explain single outcomes in macroscopic measurements, since they arguably lead to wavefunctions with macroscopically disconnected lobes, such as a dial (or “pointer”) being in one place or another or a photon being emitted in different directions, etc. Appendix B extends this reasoning to the measurement of a subsystem.
Similarly, another mechanism that could cause a loss of ergodicity is the process of copying a microscopic state onto a large number of particles (within the limits of the no-cloning theorem). This operation typically results in increasingly deeper basins in the fictitious potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$ that increase the likelihood of the particle being trapped. To see this, consider a particle in a superposition of two states $\alpha\left|1\right\rangle+\beta\left|2\right\rangle$ (with $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$) that is starting to interact with a measurement apparatus made of N1N-1 distinguishable particles, initially in a “no measurement” state $\left|0\right\rangle\otimes\cdots\otimes\left|0\right\rangle$. After a suitable unitary evolution, the joint system transitions to the state $\alpha\left|1\right\rangle\otimes\left|1\right\rangle\otimes\cdots\otimes\left|1\right\rangle+\beta\left|2\right\rangle\otimes\left|2\right\rangle\otimes\cdots\otimes\left|2\right\rangle$. Now assume that, for particle $i$, the states $\left|1\right\rangle$ and $\left|2\right\rangle$ are, respectively, associated with single-particle wavefunctions in position representation $\psi_{1}\left(x_{i}\right)=\varphi\left(x_{i}-a\right)$ and $\psi_{2}\left(x_{i}\right)=\varphi\left(x_{i}+a\right)$ for some macroscopic value $a\in\mathbb{R}^{3}$ and centrosymmetric unimodal function $\varphi\left(x_{i}\right)$ (with $x_{i}\in\mathbb{R}^{3}$). The overall wavefunction is then
Figure 2 illustrates our proposed time evolution of the wavefunction in a classic setting: the double-slit experiment. The simulation shows both the initial wave-like interference behavior and the eventual collapse to a more specific outcome as the wave packet heads towards a detector. The Supplementary materials include an animation of the same numerical experiment, as well as the simulation code itself. Appendix A also includes a treatment of a standard EPR-type experiment [23].
@@
3 Discussion
@@
The idea that the wavefunction is attracted to particles can be found in earlier work [18, 24], but our approach differs in substantial ways. Unlike [24], we show that the two-way coupling leads to Born’s rule. Unlike [18], we consider collapse towards a point in $\mathbb{R}^{3N}$ instead of a towards a single-particle density, which helps maintain EPR-type correlations between distant measurements. Also, to ensure particle trapping, our particle dynamics differs from those of both [24] and [18].
@@
The proposed theory is compatible with the standard decoherence picture (see Appendix D for a density matrix formulation): it does not affect the basic idea that interactions with the environment causes a decay of the off-diagonal terms of a subsystem’s density matrix. But once the resulting mixture of outcomes involves macroscopically distinct states, our equation of motion predicts localization towards one of these macroscopic states. Hence, it provides an objective “selection rule” that is absent under decoherence alone and avoids the need for a many-world interpretation. The decoherence concept makes it unnecessary for the wave function to collapse rapidly in our framework. The conversion from a superposition of states to a statistical mixture takes place over the same time span under our theory as under decoherence alone. The exponential decay of all but one branch of the statistical mixture, however, can take much longer: any decay rate faster than human perception (i.e., milliseconds) would be sufficient to give human observers the experience of a single measurement outcome.
The proposed theory does exhibit some undesirable features (see also Appendix C). Most obviously, the proposed modified Schrödinger equation is slightly nonlinear, whereas linearity of quantum mechanics has been verified experimentally to a great precision [25]. This places a tight upper bound on the parameter $\kappa$, thus also placing a bound on the speed at which wavefunction collapse can occur. The fact that the nonlinearity must be small mitigates other known issues associated with nonlinearity [26]. The most troubling of these issues is the possibility of faster-than-light communication. However, the objective collapse literature [27] already offers a natural solution to this possible problem that is relevant within our approach as well. Theorem 1 in [27] provides conditions under which faster-than-light communication is precluded by a suitable combination of nonlinearity with randomness. The key condition is that nonlinearity and randomness cancel out on average, so that the expected time evolution of the density matrix remains a linear operator, where the expectation is taken with respect to a random collapse field. This condition is exactly satisfied in objective collapse theories, in large part because the value of the random collapse field at any two different times are completely statistically independent. In our formalism, the role of the effective random collapse field is played by a smoothed delta function centered at the Bohmian particle’s location $r\left(t\right)$. Over a short time scale, such field unfortunately does not exhibit statistical independence. Yet, over a time scale larger than the particle’s relaxation time, the current particle’s position becomes statistically independent from the particle’s position at a given time in the past. It is known that measure-preserving dynamical systems (such as our classical particle at $r\left(t\right)$ in a potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$), typically lead to mixing processes [28]. Hence, for sufficiently large time lags, our effective random field behave analogously to that of objective collapse theories. In addition to this observation, the expected time evolution of the wavefunction under our dynamics either reduces to Schrödinger’s equation (when ergodicity holds), as shown in Equation (3), or reduces to wavefunction collapse (when ergodicity fails). This implies that the density matrix evolution remains linear and thus preclude faster-than light communication in that time regime. Within our theory, detecting faster-than light communication would entail tracking changes in the wave function over very short time scales. At the same time, if the parameter $\kappa$ is small ($\kappa\ll\sqrt{\mu^{-1}\tau}/\ell$), wavefunction evolution over short times closely follows Schrödinger’s equation and the window through which faster-than-light communication might be detected is expected to be very narrow.
@@
While our current treatment is non-relativistic, many aspects of our framework are amenable to a relativistic extension. The collapse phenomenon involves local operations: the decay term $\kappa\psi\left(x,t\right)$ is purely local while the localization terms only involves instantaneous effects over short distances (as governed by the function $\phi_{r}\left(x\right)$) and ensuring that this effect propagates at a finite speed entails no significant changes in the theory. Since relativistic versions of pilot-wave [29, 24] and objective collapse theories [30] have been developed with some success, similar ideas could carry over in the present setup. However, the presence of a global hidden variable (the particle’s position $r$) could clash with relativity, but only over time scales shorter than the particle’s relaxation time.